Which of the following ordered pairs is a solution of the system x+2y=8 and 2x-5y=-2?
Answer by ewatrrr(24785)


Hi,
No ordered pairs listed
solving the system by substitution
x+2y=8 or x = -2y + 8
2x - 5y= -2
2(-2y+8) - 5y = -2
-4y + 16 -5y = -2
18 = 9y
2 = y
solving for x
x = -2y + 8
x = -2*2 + 8
x = 4
CHECKING our Anwser
2x - 5y = -2
2*4 - 5*2 = -2